1. A car moves along a straight street. The cars initial position at time t is given by a position vector T, The cars later position vector is , at time t2 Suppose T700 mi and T - 400 mi . a. If t2-t1 = 3,0 seconds, what is the cars average speed (assuming the car only moved in one direction)? b. What was the cars average velocity? 2. A runner jogs at a stendy pace of 15 km/hr. When the runner is 7.5 km from the finish, a bird begins flying from the runner to the finish at 30 km/hr. When the bird reaches the finish line, it instantly turns around and flies back to the runner, and then turns around again, repeating the back-and-forth trips until the runner reaches the finish line. How many kilometers does the bird travel? 30 km/h finish 15 km/h .5 km
Answer #1

2. The runner goes at 15 km/hr, and has to run 7.5 km to finish. How long does that take her?

Well, 1/2 an hour (v*t = d, so t = d/v = 7.5 km / 15 km/hr = 0.5 hr) And during that half hour, the little bird is flying back and forth at a steady 30 km/hr. If it goes 30 km/hr * 1/2 an hour = 15 km, that's how far the bird went!

So the bird travels 15 km.